### v4

# ZNorm: Z-Score Gradient Normalization for Deep Neural Networks

Abstract
The rapid advancements in deep learning necessitate better training methods for deep neural networks (DNNs). As models grow in complexity, vanishing and exploding gradients impede performance. We propose Z-Score Normalization for Gradient Descent (ZNorm), an innovative technique that adjusts only the gradients to accelerate training and improve model performance. ZNorm normalizes the overall gradients, providing consistent gradient scaling across layers, thereby reducing the risks of vanishing and exploding gradients, having better performances. Our extensive experiments on CIFAR-10 and medical datasets demonstrate that ZNorm enhances performance metrics. ZNorm consistently outperforms existing methods, achieving superior results using the same experimental settings. In medical imaging applications, ZNorm improves tumor prediction and segmentation performances, underscoring its practical utility. These findings highlight ZNorm’s potential as a robust and versatile tool for enhancing the training speed and effectiveness of deep neural networks across a wide range of architectures and applications.
footnotetext:
1 Introduction
The rapid advancements in deep learning have revolutionized various fields, from computer vision to natural language processing, driven by the availability of large-scale datasets, increased computational power, and the development of sophisticated network architectures. However, training deep neural networks (DNNs) [9] remains a formidable challenge. As models become deeper and more complex, issues such as vanishing and exploding gradients frequently arise, impeding model performance [11].
To address these challenges, several normalization techniques have been proposed to improve the training process. Methods such as batch normalization (BN) [5], layer normalization (LN) [1], and weight standardization (WS) [15] have been widely adopted for their effectiveness in stabilizing training by adjusting the mean and variance of activations or weights. Another promising recent approach, Gradient Centralization [22], focuses on normalizing the gradients themselves by adjusting their mean to zero, which has shown to improve training consistency and performance. Similarly, Gradient Clipping [12] has been employed to prevent the exploding gradient problem by capping gradients at a predefined threshold. Weight decay [8], another widely used technique, applies L2 regularization to the weights during training to prevent overfitting and control weight growth, which indirectly improves the stability of the training process.
Building on these approaches, our work aims to demonstrate that by simply normalizing and adjusting gradients, significant improvements in performance can be achieved across a variety of network architectures. We introduce ZNorm, a technique designed to normalize gradients, providing consistent gradient scaling. This enhances the generalization capabilities of DNNs. Unlike other methods that primarily focus on activation or weight normalization, ZNorm directly addresses the gradients, leading to faster training. Instead of striving to develop increasingly complex architectures, our approach suggests that substantial performance gains can be realized by refining the training process itself. By improving how gradients are handled during training, we achieve better results with existing architectures, thereby challenging the notion that more complex models are always necessary for better performance.
Figure 1: Visualization of the ZNorm($\Phi$) process applied to gradient matrices and tensors in both fully-connected and convolutional layers. The process includes calculating the gradient mean vector and standard deviation vector, followed by normalizing the gradients via ZNorm.
The impact of ZNorm is particularly noteworthy as it achieves these performance gains, making it an accelerator for improving training processes. Our extensive experiments on both synthetic and real-world datasets validate the effectiveness of ZNorm. The results consistently show that ZNorm boosts performance metrics, leading to more accurate models. For instance, in medical imaging applications, ZNorm has demonstrated its practical utility by improving the accuracy of tumor prediction and segmentation tasks. This makes ZNorm a valuable tool for advancing deep learning models across a wide range of architectures and applications by focusing on the crucial aspect of gradient normalization.
2 Related Works
Optimization techniques are central to training deep neural networks effectively. Stochastic gradient descent (SGD) and its variants, such as SGD with momentum [14] and adaptive methods like Adam [6], have been crucial in optimizing complex architectures. These methods aim to minimize the loss function via parameter updates. Adaptive optimizers like Adagrad [2] and RMSProp [19] further adjust learning rates dynamically based on gradient magnitude and history, stabilizing the training process for deep models.
Normalization techniques help mitigate training instability and smooth out optimization landscapes. Batch normalization (BN) [5] normalizes activations across mini-batches to reduce internal covariate shifts. Layer normalization (LN) [1] and group normalization (GN) [21] normalize along different dimensions, such as layers or groups of channels, while weight normalization (WN) [17] and weight standardization (WS) [15] ensure consistent weight scaling, contributing to faster and more stable training.
Gradient adjustment methods directly manipulate the gradient updates to improve training stability and prevent issues like exploding or vanishing gradients. Gradient Clipping [13] is used to prevent the exploding gradient problem by capping the gradient magnitudes. If the gradient exceeds a predefined threshold $\tau$, it is rescaled like $\nabla_{\theta}L(\theta_{t})\leftarrow\nabla_{\theta}L(\theta_{t})\cdot\min\left(1,\frac{\tau}{||\nabla_{\theta}L(\theta_{t})||}\right)$. This ensures that no gradient update is too large, helping stabilize training in recurrent or deep networks.
AdamW [10] improves upon traditional weight decay by decoupling L2 regularization from the gradient update in adaptive optimization methods like Adam. In traditional Adam, weight decay is applied implicitly through the L2 regularization term, which affects both the gradient and the parameter updates. AdamW, however, applies weight decay directly to the weights during the parameter update step, ensuring that the regularization only affects the weight magnitudes and not the gradient direction. The gradient update for weight decay in AdamW is modified as $\nabla_{\theta}L_{\text{decay}}(\theta_{t})=\nabla_{\theta}L(\theta_{t})+\lambda\theta_{t}$ where $\lambda$ is the weight decay coefficient. This decoupled approach in AdamW prevents the regularization term from interfering with the adaptive learning rate updates, leading to more stable training and better generalization, especially in large-scale models. Gradient Centralization [22] normalizes the gradients by subtracting their mean, ensuring a zero-mean gradient. The gradient update is adjusted as $\nabla_{\theta}L(\theta_{t})\leftarrow\nabla_{\theta}L(\theta_{t})-\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta_{i}}L(\theta_{t})$ where $n$ is the number of gradient elements. This technique improves generalization and speeds up convergence by regularizing the training dynamics.
Our proposed Z-Score Normalization (ZNorm) builds on these techniques. By applying Z-score normalization directly to overall gradients, we standardize gradient scaling across layers, reducing vanishing and exploding gradients [12] while ensuring consistent learning progress. This leads to improved generalization and robustness of the models.
3 Methodology
In this section we introduces Z-Score Normalization (ZNorm), a technique designed to standardize gradient tensors across all network layers.
Algorithm 1 Adam with Z-Score Normalization (ZNorm)
Input: Weight vector $\theta^{0}$, step size $\mu$, $\beta_{1}$, $\beta_{2}$, $\epsilon$, $\mathbf{m}^{0}$, $\mathbf{v}^{0}$
Training step:
1:for $t=1$ to $T$ do
2:     $\mathbf{g}^{t}=\nabla\mathcal{L}(\mathbf{\theta})$triangleright Compute gradient
3:     $\hat{\mathbf{g}}^{t}=\Phi_{ZNorm}(\mathbf{g}^{t})$triangleright Apply Z-Score Normalization
4:     $\mathbf{m}^{t}=\beta_{1}\mathbf{m}^{t-1}+(1-\beta_{1})\hat{\mathbf{g}}^{t}$
5:     $\mathbf{v}^{t}=\beta_{2}\mathbf{v}^{t-1}+(1-\beta_{2})\hat{\mathbf{g}}^{t}\odot\hat{\mathbf{g}}^{t}$
6:     $\hat{\mathbf{m}}^{t}=\frac{\mathbf{m}^{t}}{1-\beta_{1}^{t}}$
7:     $\hat{\mathbf{v}}^{t}=\frac{\mathbf{v}^{t}}{1-\beta_{2}^{t}}$
8:     $\mathbf{\theta}^{t+1}=\mathbf{\theta}^{t}-\mu\frac{\hat{\mathbf{m}}^{t}}{\sqrt{\hat{\mathbf{v}}^{t}}+\epsilon}$
9:end for
3.1 Notations
Consider a deep neural network consisting of $L$ layers, where each layer $l$ has associated weights $\mathbf{\theta}^{(l)}$. Fully connected layers can be expressed as $\mathbf{\theta}^{(l)}_{fc}\in\mathbb{R}^{D_{l}\times M_{l}}$, where $D_{l}$ represents the number of neurons and $M_{l}$ represents the dimension of the input to the layer. Convolutional layers, $\mathbf{\theta}^{(l)}_{conv}$, can be represented as a 4-dimensional tensor $\mathbf{\theta}^{(l)}_{conv}\in\mathbb{R}^{C_{out}^{(l)}\times C_{in}^{(l)}\times k_{1}^{(l)}\times k_{2}^{(l)}}$, where $C_{in}^{(l)}$ and $C_{out}^{(l)}$ are the number of input and output channels, respectively, and $k_{1}^{(l)}$ and $k_{2}^{(l)}$ represent the kernel sizes. Let $\nabla\mathcal{L}(\mathbf{\theta}^{(l)})$ denote the gradient of the loss function $\mathcal{L}$ with respect to the weights $\mathbf{\theta}^{(l)}$ of layer $l$. $\nabla\mathcal{L}(\mathbf{\theta})$ denotes the overall gradient tensor.
3.2 Z-Score Normalization for Gradients
Z-Score Normalization is applied layer-wise, meaning each layer’s gradient $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ is individually normalized using ZNorm, which can be expressed as:
$\Phi_{ZNorm}(\nabla\mathcal{L}(\mathbf{\theta_{(t)}^{(l)}}))=\frac{\nabla\mathcal{L}(\mathbf{\theta_{(t)}^{(l)}})-\mu_{\nabla\mathcal{L}(\mathbf{\theta_{(t)}^{(l)}})}}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta_{(t)}^{(l)}})}+\epsilon}$ (1)
Here, $\mu_{\nabla\mathcal{L}(\mathbf{\theta^{l}})}$ represents the mean of the gradients in layer $l$, and $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{l}})}$ is their standard deviation. A small constant $\epsilon$ (typically 1e-10) is added to prevent division by zero. When we refer to $\Phi_{ZNorm}(\nabla\mathcal{L}(\mathbf{\theta}))$, it denotes the collection of all layer-wise gradients that have been individually normalized by ZNorm across the entire network. ZNorm automatically scales the gradients within each layer without requiring additional hyperparameters.
3.3 Embedding ZNorm to Adam
Z-Score Normalization (ZNorm) can be seamlessly integrated into existing deep neural network (DNN) optimization algorithm, Adam [6]. ZNorm, applied directly to the gradient tensors. In this paper, we specifically focus on embedding ZNorm into the Adam optimization algorithm, which is one of the most widely used optimization algorithms in deep learning due to its adaptive learning rate properties and effectiveness in training deep models. Algorithm 1 shows how ZNorm can be embedded into the Adam algorithm with minimal changes. The ZNorm step is applied immediately after computing the gradient, and then the normalized gradient is utilized in the subsequent Adam updates.
4 Theoretical Analysis
Z-score normalization normalizes gradients by subtracting their mean and dividing by the standard deviation, potentially altering the direction of individual components. This raises concerns about whether the overall descent direction is maintained. This analysis explains why, despite these changes, the optimization process remains effective. Specifically, it shows that the inner product of the normalized gradient with the original gradient remains positive, preserving the overall descent direction necessary for reducing the objective function.
Lemma 4.1. Let $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\in\mathbb{R}^{d}$ be the original gradient vector with mean $\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}=\frac{1}{d}\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{i}^{(l)}})$, variance $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}={\frac{1}{d}\sum_{i=1}^{d}(\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})})^{2}}$. The Z-score normalized gradient $\Phi_{Z}(\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}}))$ is defined as
$\Phi_{Z}(\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}}))=\frac{\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}$
for $\text{for }i=1,2,\dots,d.$
Then, the inner product of the normalized gradient $\Phi_{Z}$ with the original gradient $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ is given by
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\frac{\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}-d\mu_{\nabla}\mathcal{L}(\mathbf{\theta^{(l)}})^{2}}{\sigma_{\nabla}\mathcal{L}(\mathbf{\theta^{(l)}})},$
where $\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}$ is the Euclidean norm of the vector $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$. This result indicates that while the Z-score standardization can alter the direction of individual components of the gradient, the overall descent direction is preserved.
Proof. We begin by recalling the definition of the Z-score standardized gradient:
$\Phi_{Z}(\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}}))=\frac{\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}.$
where $\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}=\frac{1}{d}\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})\quad\text{and}\quad\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}=\frac{1}{d}\sum_{i=1}^{d}(\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})})^{2}$ Here, $\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}$ represents the mean of the gradient components, and $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}$ represents the standard deviation, which measures the spread of the gradient components around the mean.
Next, the z-score normalized gradient vector $\Phi_{Z}$ can be expressed in vector form as:
$\Phi_{Z}=\frac{1}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}(\nabla\mathcal{L}(\mathbf{\theta^{(l)}})-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}\mathbf{1}),$
where $\mathbf{1}$ is the $d$-dimensional vector of ones.
Here, the inner product $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, which is given by:
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\left(\frac{(\nabla\mathcal{L}(\mathbf{\theta^{(l)}})-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}\mathbf{1})}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}\right)^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}}).$
Expanding this expression, we get:
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ $=\frac{((\nabla\mathcal{L}(\mathbf{\theta^{(l)}}))^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}$ (2)
$-\frac{-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}\mathbf{1}^{T}{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})})}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}$ (3)
Now, we calculate $\mathbf{1}^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, which represents the sum of all components of the gradient vector $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$:
$\mathbf{1}^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})=d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}.$
Substituting this into the expression for $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, we obtain:
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\frac{(\nabla\mathcal{L}(\mathbf{\theta^{(l)}}))^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})-d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}.$
Next, let’s express $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ in terms of the Euclidean norm of $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$:
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\sum_{i=1}^{d}(\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}}))^{2}=\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}.$
Thus, the inner product becomes:
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\frac{\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}-d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}.$
This equation shows that the inner product of the Z-score normalized gradient $\Phi_{Z}$ with the original gradient $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ is determined by the difference between the squared Euclidean norm of $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ and the scaled mean squared value of $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, divided by the standard deviation $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}$.
This result is significant because it indicates that while the Z-score standardization can change the direction of individual components of the gradient, the overall directionality, as measured by the inner product $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, remains aligned with the original gradient direction. Specifically, if $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ is a descent direction, then $(\Phi_{Z})^{T}$ will also be a descent direction, as the inner product $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ remains positive, which proves the lemma.
Theorem 4.2. Given the same conditions as in the Lemma, the Z-score normalized gradient $\hat{g}$ preserves the descent direction of the original gradient $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$. Specifically, if $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ is a descent direction, then $(\Phi_{Z})$ is also a descent direction, i.e.,
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})>0.$
This implies that even though Z-score standardization may change the directions of individual gradient components, the overall gradient vector still points in a direction that will reduce the objective function.
Proof. From the result of the Lemma, we have:
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\frac{\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}-d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}.$
Here, $\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}=\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})^{2}$ is the squared Euclidean norm of the gradient vector $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, and $\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}$ is the mean of the gradient components.
Since $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}$ is the standard deviation of the elements of $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, we have $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}>0$. The standard deviation $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}$ is positive as long as the components of $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ are not all identical (i.e., $g_{i}$ are not constant for all $i$).
Next, consider the term $d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}$. This term can be rewritten as:
$d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}=d\left(\frac{1}{d}\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})\right)^{2}=\frac{1}{d}\left(\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})\right)^{2}.$
This expression represents the squared mean of the gradient components scaled by the dimension $d$. Now, by the Cauchy-Schwarz inequality, we have:
$\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})^{2}\geq\frac{1}{d}\left(\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})\right)^{2}=d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}.$
This inequality ensures that the term $\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}-d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}$ is non-negative:
$\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}-d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}\geq 0.$
Therefore, the inner product $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ can be expressed as:
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\frac{\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}-d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}\geq 0.$
For $\hat{g}^{T}g$ to be strictly positive, $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ must not be a constant vector (i.e., not all $\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})$ are the same), ensuring that $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}>0$ and $\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}>d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}$. This condition holds under typical scenarios in optimization where the gradient vector $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ varies across its components. The above proof shows that $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ is positive under general conditions, but it may not always be positive in specific cases. Particularly, if all components of the gradient are identical (i.e., $\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})=\text{constant}$), the standard deviation $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}=0$, making the inner product undefined or zero. Therefore, while Z-score standardization generally preserves the descent direction in typical scenarios, it may not always do so in extreme cases. Therefore, under these conditions, $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})>0$, which implies that the Z-score normalized gradient $(\Phi_{Z})$ indeed preserves the descent directionality. Despite the potential changes in individual components caused by Z-score standardization, the overall direction of the gradient vector $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ that leads to a decrease in the objective function is preserved, proving the theorem.
5 Experimental Results
In this section, we conducted a series of experiments on both image classification and image segmentation tasks to show the performances of our Z-Score Gradient Normalization (ZNorm) method.
5.1 Experimental Setting
All experiments were performed using the Adam optimizer [6], with the baseline referring to the standard Adam gradient. The experiments involved Gradient Clipping [13], Gradient Centralization [22], and ZNorm (Ours). Additionally, Weight Decay 3 and Weight Decay 4 correspond to decay rates of 1E-3 and 1E-4, respectively. For clipping, the value was 0.1. For image classification, we used the CIFAR-10 [7] and PatchCamelyon [20] datasets. All datasets were not augmented.
Both were trained with a batch size of 256, a learning rate of 0.001, and 100 epochs for CIFAR-10 and 50 epochs for PatchCamelyon. For image segmentation, we used the LGG MRI dataset [23] for brain tumor segmentation. Training settings were a batch size of 128, an initial learning rate of 0.01, and total training for 50 epochs, with the learning rate reduced by a factor of 10 every 5 epochs starting from epoch 30.
Table 1: Performance Comparison on CIFAR-10 Dataset [7] for various convolutional neural network architectures and gradient normalization techniques. The table shows the test accuracy and train loss across several models and normalization methods. Bold values represent the highest test accuracy achieved for each model. ZNorm consistently outperforms other methods in terms of test accuracy, demonstrating its effectiveness in improving model performance.
[…]
5.2 Image Classification Results for CIFAR-10
In this section, we assess the effectiveness of ZNorm on the CIFAR-10 dataset [7] across various neural network architectures. ZNorm consistently improves performance across different CNN architectures, offering better test accuracy and stability compared to other methods.
Table 1 provides detailed experimental results, showing how ZNorm enhances performance across multiple models. For instance, in ResNet-56 [3], ZNorm achieves a test accuracy of 0.812, outperforming the baseline (0.802) as well as methods like Gradient Centralization (0.804) and Gradient Clipping (0.747). Similarly, ResNet-101 [3] shows improved results with ZNorm, reaching a test accuracy of 0.820, surpassing the baseline (0.770) and other techniques.
In DenseNet-121 [4], ZNorm boosts test accuracy to 0.799, outperforming the baseline (0.759) and Gradient Centralization (0.784). DenseNet-169 [4] also benefits from ZNorm, achieving the highest test accuracy of 0.802, compared to the baseline (0.766) and other methods.
Finally, MobileNetV2 [18] with ZNorm reaches a test accuracy of 0.7801, outperforming the baseline (0.763) and other techniques. These results demonstrate ZNorm’s consistent ability to improve accuracy across a diverse set of models and architectures.
Table 2: Performance comparison of Breast Tumor Prediction on PatchCamelyon [20] dataset based on ResNet [3] models with different normalization techniques. Bold values represent the highest test accuracy achieved for each model. ZNorm consistently outperforms other methods in terms of test accuracy, demonstrating its effectiveness in improving model performance
[…]
Table 3: Performance Comparison of Brain Tumor Segmentation on LGG MRI Dataset [23] with various architectures and normalization methods. Bold values represent the highest test accuracy achieved for each model. ZNorm consistently outperforms other methods in terms of Test Tversky, demonstrating its effectiveness in improving model performance
[…]
5.3 Image Classification Results for Breast Tumor
We evaluated the effectiveness of the Z-Score Normalization (ZNorm) method on the PatchCamelyon dataset [20], which consists of breast cancer pathology images used to determine whether the patchified images contain tumors. This real-world dataset allows us to further assess the performance of ZNorm beyond the controlled CIFAR-10 experiments.
We experimented with ResNet-56, ResNet-101, and ResNet-152 models [3]. As shown in Table 2, ZNorm consistently outperformed other normalization techniques across these models. For ResNet-56, ZNorm achieved a test accuracy of 0.915, significantly higher than the baseline accuracy of 0.880 and other methods such as Gradient Centralization (0.887) and Gradient Clipping (0.887). Similarly, for ResNet-101, ZNorm reached the highest test accuracy of 0.917, surpassing both the baseline (0.870) and the other methods. In the case of ResNet-152, ZNorm also outperformed other approaches, achieving a test accuracy of 0.912, compared to the baseline accuracy of 0.884.
These results demonstrate that ZNorm can effectively improve performance on real-world datasets like PatchCamelyon, validating its potential in medical imaging tasks such as tumor prediction.
Figure 2: Comparison of segmentation mask results using different methods such as GC[22], Clipping[13], Weight Decays [10] and ZNorm on LGG datasets[23] based on ResNet-50-Unet [16]. ZNorm demonstrates superior performance, producing segmentation masks that more closely match the ground-truth compared to other methods.
5.4 Image Segmentation for Brain Tumor
For the brain tumor segmentation task, we utilized the LGG MRI dataset [23], which is aimed at detecting and segmenting tumors in brain MRI images. Three well-established models in the domain were applied: ResNet50-Unet [16], U-net++ [24], and Attention UNet [25].The performance comparison shown in Table 3 highlights that ZNorm consistently outperforms other methods across various models. In the ResNet50-Unet [16], ZNorm achieved the highest test F1 score of 0.917, a Tversky accuracy of 0.914, and the lowest Hausdorff distance of 2.663, outperforming the baseline and other gradient handling techniques like Gradient Centralization and Weight Decay. Similarly, U-net++ [24] produced a strong Tversky accuracy of 0.910 with ZNorm, coupled with competitive results in other metrics, further demonstrating its effectiveness. In the Attention UNet [25], ZNorm recorded the best Tversky accuracy (0.928) and the lowest Hausdorff distance (2.947), while maintaining a high F1 score of 0.894. These findings demonstrate that ZNorm consistently delivers more accurate and precise tumor segmentation compared to other normalization techniques.
The qualitative results in Figure 2 illustrate how ZNorm produces segmentation masks that closely match the ground-truth, with significant improvements over baseline methods, Gradient Centralization [22], Gradient Clipping [13], and different weight decay settings [10]. These results clearly show that ZNorm enhances the robustness and accuracy of brain tumor segmentation across multiple architectures.
6 Conclusion
We proposed ZNorm, a novel method for gradient normalization, which has proven to significantly enhance the performance of deep neural networks. By normalizing both the mean and variance of the gradients, ZNorm offers consistent gradient scaling across layers, effectively mitigating issues like vanishing and exploding gradients. Our experimental results, spanning across a range of architectures and tasks such as image classification on CIFAR-10 and PatchCamelyon datasets and brain tumor segmentation on the LGG MRI dataset, consistently demonstrate that ZNorm outperforms existing methods like Gradient Centralization and Gradient Clipping. ZNorm leads to better generalization and higher accuracy in both classification and segmentation tasks compared to other normalization techniques. These results highlight ZNorm’s versatility and robustness, making it a valuable tool for improving the performance of deep learning models across various applications and architectures. Given these advantages, ZNorm stands out as a promising method for future work in deep neural network optimization and beyond.